OpenStudy (anonymous):

Solve y – 3x = 5 by the substitution method. y + x = 3

OpenStudy (anonymous):

@ParthKohli , by substitution not elimination

Parth (parthkohli):

My mistake. Subtracting \(x\) from the second equation fetches ya \( \color{Black}{\Rightarrow y = 3 - x}\) Now substitute that back into the first equation: \( \color{Black}{\Rightarrow (3 - x) - 3x = 5}\)

OpenStudy (anonymous):

@ParthKohli now you got it

Parth (parthkohli):

Yes. I sometimes fail absent-mindedly.

OpenStudy (anonymous):

ok..?

OpenStudy (anonymous):

@coolaidd you got it?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3 - x - 3x = 5 }\) \( \color{Black}{\Rightarrow 3 - 4x = 5}\) I'd like coolaid to do the rest.

Parth (parthkohli):

My weakness is that I am a lazy person, so I am not going to do everything else.

hero (hero):

I personally think @coolaidd should be trying some of these on her own.

OpenStudy (anonymous):

y-3x=5 y+x = 3 We have to put y= ..... or x = ..... on one side of both equations, hence the easiest way is to put y on side since the first equation you can cancel the x on the left and add it on the right. Here is our two new equations y=3x+5 y= 3-x Since both equations are equivalent to y, we can make the equivalent for the same equation. (Basically we turn both into one single equation..) 3x+5= 3-x => 3x+x = 3-5 => 4x = -2 => x = -1.5 Now subsitute -1.5 in the second original equation (since its the easiest) and now y=4.5 That's how you do it!

OpenStudy (anonymous):

sorry x = -0.5 and y = 3.5, sorry my mistake! that's the answer

OpenStudy (anonymous):

You can also do @ParthKohli did, but it doesn't really matter.

OpenStudy (anonymous):

@vmathhelp but its good to follow the questions instructions